OpenStudy (anonymous):

Which expression is equivalent to 64  x2?

OpenStudy (anonymous):

please rewrite the question

OpenStudy (anonymous):

can you tell me how to sovle this problem step by step please which expression is equivalent to 64 x2

OpenStudy (anonymous):

\[64- x2\]

OpenStudy (anonymous):

what expressions are there? Is this a multiple answer question?

OpenStudy (anonymous):

this is a regents question

OpenStudy (anonymous):

Which expression is equivalent to 64 x2? (1) (8 x)(8 x) (3) (x 8)(x 8) (2) (8 x)(8 x) (4) (x 8)(x 8)

OpenStudy (anonymous):

64 - x^2. this is difference of two squares. (a^2 - b^2) = (a+b)(a-b)

OpenStudy (anonymous):

so what would the answer be to my question i want to now how u get the answer step by step

OpenStudy (anonymous):

what is the square root of 64?

OpenStudy (anonymous):

oo okay you have to find a number that mulitplys in to 64 equally

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

If you want to know, a difference in squares pattern works when you find the square root of both terms and have two binomials, one with a + and one with a -. (x + 3)(x - 3) = x^2 + 3x - 3x - 9 = x^2 - 9 See how the middle terms cancel out? Your problem is the same, except that the x is the second term.